Two angles are complementary. The sum of the measure of the first angle and half of the second angle is 72.5*. Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?

Answer by nerdybill(7384)

Two angles are complementary. The sum of the measure of the first angle and half of the second angle is 72.5*. Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?
.
If two angles are complementary, the sum of the two angles equals 90 degrees.
.
Let x = first angle
then
90-x = second angle
.
From:"The sum of the measure of the first angle and half of the second angle is 72.5" we get our equation:
x + (1/2)(90-x) = 72.5
2x + (90-x) = 145
x + 90 = 145
x = 55 degrees (first angle - other angle)
.
Second angle:
90-x = 90-55 = 35 degrees (second angle-smaller angle)